The sum of three consecutive numbers is twelve less than six times the lowest number. What are the three numbers?
Answer by richwmiller(17219)

x+x+1+x+2=6x-12
3x+3=6x-12
15=3x
x=5
x+1=6
x+2=7
check
5+6+7=6*5-12
18=30-12
18=18
ok